OpenStudy (anonymous):

Explain how you could use the Pythagorean Theorem to find the distance between the points the coordinate pair zero comma negative three and the coordinate pair three comma four. Find the distance. You can type carat two to indicate "squared" and "sqrt" to indicate "square root." (Examples: five carat two equals five squared and sqrtsqrt twenty-five equals the square root of twenty-five)

OpenStudy (mathstudent55):

Do you still need help with this problem?

OpenStudy (anonymous):

d^2=x^2=Y^2, d^2=(x2-x1)^2+(y2-Y1)^2 message me if this is what you were looking for